OpenStudy (anonymous):

explain this rule for exponents: x^(-n) = 1/x^n

OpenStudy (anonymous):

what does "explain" mean in this context?

OpenStudy (anonymous):

if it means "give an example" you can say \[10^{-2}=\frac{1}{10^2}=\frac{1}{10}=0.01\]

OpenStudy (anonymous):

if it means 'why is it true?' the answer would be "it is true by definition"

OpenStudy (anonymous):

it says in my own words, explain this rule for exponents

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

then use your own words

OpenStudy (anonymous):

^ I was going to say that, but I was afraid it'd be too cheeky ^.^

OpenStudy (anonymous):

@PeterPan new around here?

OpenStudy (anonymous):

Ten days, or eleven.

OpenStudy (anonymous):

the cheek will come take my word for it

OpenStudy (anonymous):

I would but I don't understand what they are wanting

OpenStudy (anonymous):

@HelenKizer really use your own words what it says is exactly what it says, that \(x^{-n}=\frac{1}{x^n}\) i.e. to take something to a power that is negative, it means take the reciprocal of that number raised to the positive power

OpenStudy (anonymous):

for example \[2^{-3}=\frac{1}{2^3}=\frac{1}{8}\] or \[\left(\frac{1}{4}\right)^{-3}=4^3=64\] or \[\frac{1}{5^{-2}}=5^2=25\]

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

Or you can sound fancy, and write this down... \[\huge 1=x^0=x^{n-n}=x^{-n}x^n\]

OpenStudy (anonymous):

\[\Huge \frac1{x^n}=x^{n-1}\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

oops? :/

OpenStudy (anonymous):

typo there